OpenStudy (anonymous):

Find the integral. (5x^(-3) - 4x^(-1))

OpenStudy (anonymous):

\[(5x^{-3}-4x ^{-1})\] This is the equation.

OpenStudy (blockcolder):

Use the ff. formulas: \[\int x^n\ \text{d}x=\frac{x^{n+1}}{n+1}+C \text{ for }x\neq-1\\ \int x^{-1}\ \text{d}x=\ln|x|+K\] where C and K are arbitrary constants.

OpenStudy (anonymous):

So would it be \[-5/2x^{-2}-4\ln \left| x \right|\]? I get really confused with the natural log when finding the integral.

OpenStudy (blockcolder):

Yep. That's right.

OpenStudy (anonymous):

Awesome! Your a math genius! Thanks again!

OpenStudy (blockcolder):

And don't forgot the constant btw. :D